If you buy a home with less than 20% down, you will pay an additional monthly fee, PMI (private mortgage insurance), until you reach 80% equity. Keep track of when you reach 80% equity so you can request to have your PMI removed. Ken Buckmiller’s home recently appraised at $290,000. His mortgage was for $275,000 at 5% for 30 years with PMI of $229.17 per month. What is his monthly payment plus PMI? His mortgage balance is currently $222,990. Has he reached 80% equity?
Answer #1

Monthly payment: PV = -275,000; N = 30*12 = 360; rate = 5%/12 = 0.4167%, solve for PMT.

PMT = 1,476.26

PMI = 229.17

Total monthly payment = 1,476.26 + 229.17 = 1,705.43

For 80% equity, mortgage balance has to be 80%*290,000 = 232,000

His mortgage balance is 222,990 so he has reached 80% equity.